### v1

# EMFusion: Conditional Diffusion Framework for Trustworthy Frequency Selective EMF Forecasting in Wireless Networks

Abstract
The rapid growth in wireless infrastructure has increased the need to accurately estimate and forecast electromagnetic field (EMF) levels to ensure ongoing compliance, assess potential health impacts, and support efficient network planning. While existing studies rely on univariate forecasting of wideband aggregate EMF data, frequency-selective multivariate forecasting is needed to capture the inter-operator and inter-frequency variations essential for proactive network planning. To this end, this paper introduces EMFusion, a conditional multivariate diffusion-based probabilistic forecasting framework that integrates diverse contextual factors (e.g., time of day, season, and holidays) while providing explicit uncertainty estimates. The proposed architecture features a residual U-Net backbone enhanced by a cross-attention mechanism that dynamically integrates external conditions to guide the generation process. Furthermore, EMFusion integrates an imputation-based sampling strategy that treats forecasting as a structural inpainting task, ensuring temporal coherence even with irregular measurements. Unlike standard point forecasters, EMFusion generates calibrated probabilistic prediction intervals directly from the learned conditional distribution, providing explicit uncertainty quantification essential for trustworthy decision-making. Numerical experiments conducted on frequency-selective EMF datasets demonstrate that EMFusion with the contextual information of working hours outperforms the baseline models with or without conditions. The EMFusion outperforms the best baseline by 23.85% in continuous ranked probability score (CRPS), 13.93% in normalized root mean square error, and reduces prediction CRPS error by 22.47%.
I Introduction
The rapid proliferation of new wireless communication technologies, such as transmissions at higher frequencies, massive antennas, and reconfigurable intelligent surfaces, has led to an unprecedented increase in the radio-frequency (RF)-based electromagnetic field (EMF) sources [mootoo2025emforecaster, chiaraviglio2021health]. This technological surge has amplified public and regulatory concerns regarding potential human health effects associated with EMF exposure [chiaraviglio2021health, icnirp2020guidelines]. Consequently, accurate EMF monitoring and forecasting is becoming indispensable for regulatory compliance and for optimizing network deployment and resource management in a proactive manner [mootoo2025emforecaster, stam2018comparison].
To date, existing research studies in the domain focused on univariate forecasting of aggregate EMF exposure collected across a large bandwidth [mootoo2025emforecaster]. However, knowing the wideband EMF exposure is often not enough for active network planning or proactive network resource allocation. Instead a detailed frequency selective EMF exposure breakdown is necessary to identify dominant exposure sources and pin-point the EMF contributions from specific technologies (e.g., 3G, 4G, and 5G deployments) and network operators. Frequency selective EMF forecasting enables regulatory compliance with higher granularity, optimize network deployment as well as spectrum and power allocation policies. Nevertheless, due to the scarcity of frequency selective EMF datasets, forecasting EMF contributions at specific transmission frequencies with multiple network operators and/or diverse wireless technologies remain a fundamental challenge.
Along another note, the uncertainty in EMF patterns, driven by fluctuating network traffic, user mobility, and propagation conditions, means that deterministic point forecasts are insufficient due to the significant errors they produce. Probabilistic forecasting approaches can provide a comprehensive understanding of potential EMF exposure levels, including their likelihood, range, and associated uncertainties [huang2025probabilistic].
Over the past decade, deep learning (DL) solutions have achieved remarkable progress in time-series forecasting (TSF), yet the trained models remain limited in their ability to represent uncertainty and complex temporal dynamics. Conventional architectures such as recurrent neural networks (RNNs) and transformers are primarily designed for deterministic prediction, producing single-point forecasts that often overlook the inherent randomness of real-world systems. As a result, these models can yield overconfident estimates and struggle to generalize under noisy, incomplete, or dynamic conditions. Moreover, their autoregressive nature can lead to error accumulation over extended horizons, reducing long-term reliability. Recently, Generative diffusion models (DMs) enable learning the full probability distribution of future trajectories through an iterative denoising process [huang2025probabilistic, sohl2015deep, ho2020denoising]. This approach enables them to generate diverse and realistic forecasts, impute missing or corrupted data, transmit semantic wireless signals with high fidelity [ref:imputation, 10896580], and flexibly incorporate contextual variables such as seasonal, holiday, or working hour trends [borovykh2017conditional]. By combining probabilistic reasoning with stable and interpretable generation, diffusion-based forecasting represents a significant step forward in capturing the uncertainty and complexity of temporal phenomena.
TABLE I: Classification of deep learning models for wireless networks forecasting/prediction applications. UV: Univariate forecasting; MV: Multivariate forecasting; Cond.: External Conditions; Imp.: Imputation; Trust.: Trustworthiness.
Domain Ref. Model Target Dataset UV/MV Cond. Imp. Trust.
Network Traffic [nikravesh2016mobile] MLP, MLPWD, SVM Commercial LTE trial traffic UV/MV
[di2023multivariate] VAR, RNN LTE-A VoIP traffic UV/MV
[dalgkitsis2018traffic] LSTM RNN Vodafone 4G throughput UV
[habib2024transformer] Transformer 5G/LTE O-RAN aggregate traffic UV
[fiandrino2024aichronolens] LSTM + XAI 4G load and active-user datasets UV
[isravel2024multivariate] VAR, LSTM ubiquitous healthcare traffic MV
Channel Prediction [sone2020wireless] Holt–Winters, LSTM Real enterprise WLAN traffic UV/MV BS Traffic
[zhang2021deep] LSTM (SCP), RNN Synthetic LEO mMIMO CSI UV
QoS Prediction [hameed2022toward] Temporal Transformer Real IoT smart-building datasets UV/MV
[dinaki2021forecasting] BiLSTM–CNN CDN video streaming QoE dataset MV
[colpitts2023short] LSTM, Seq2Seq Rural fixed-wireless LTE KPIs MV Weather
[zhang2025probabilistic] Diffusion + TCN Real DTN-inspired sensor datasets MV Tempreature
EMF Exposure [kiouvrekis2024comparative] k-NN, XGBoost EMF measurements UV Population Density XAI
[bakcan2022measurement] CNN Wideband RF-EMF UV
[pala2021examining] LSTM, RNN Wideband RF-EMF UV
[nguyen2024deep] Transformer, CNN Wideband RF-EMF MV
[mootoo2025emforecaster] Patching, MLP Wideband RF-EMF UV
EMFusion CDM + Cross attention Frequency Selective RF-EMF UV/MV Time, Date, Season
To this end, this paper proposes EMFusion, a conditional diffusion model (CDM) for multivariate probabilistic forecasting of EMF exposure at different frequencies. The considered frequency-selective EMF datasets are multivariate since each timestamp contains a set of EMF measurements across multiple frequency bands, wireless technologies, and network operators, forming a multidimensional observation vector. Modeling these jointly, rather than independently, enables capturing interdependencies such as shared temporal patterns, correlated behavior, and coupled exposure dynamics. As such, EMFusion is capable to model the joint distribution of multi-frequency EMF exposure trajectories, capture complex inter-frequency dependencies, and provide built-in uncertainty quantification directly from the generative process, eliminating the need for post-hoc calibration wrappers. To our best knowledge, none of the existing research in the domain of wireless networks have applied conditional DMs for TSF in general and for EMF TSF in particular. The specific contributions of this paper can thus be summarized as follows:
  1. 1.
    The proposed EMFusion is based on a CDM for multivariate probabilistic forecasting of frequency-specific EMF exposure that allows the generation process to be guided by various contextual factors or conditions. EMFusion leverages contextual variables such as time of the day, day of the week, holiday/working days, working hours, and seasonal trends. This conditioning not only enhances the accuracy and relevance of the forecasts, but also allows for the generation of scenarios tailored to specific environmental states. The proposed architecture features a residual U-Net backbone enhanced by a cross-attention mechanism that dynamically integrates external conditions to guide the generation process. Furthermore, EMFusion integrates an imputation-based sampling strategy that treats forecasting as a structural inpainting task, ensuring temporal coherence even with irregular or missing EMF measurements.
  2. 2.
    Unlike conventional point forecasters that output a single trajectory, we propose a robust interval estimation framework based on the generative capabilities of EMFusion. By treating the ensemble of generated trajectories as samples from a conditional distribution, we employ Kernel Density Estimation (KDE) to reconstruct continuous probability density functions at each time step. This approach enables the derivation of empirically calibrated prediction intervals without imposing parametric Gaussian assumptions, thereby achieving empirically validated coverage at prescribed confidence levels and providing trustworthy uncertainty quantification on EMF multivariate forecasting rather than simple point estimation.
  3. 3.
    EMFusion can be customized for both univariate and multivariate forecasting scenarios. Multivariate forecasting captures correlation across distinct network operators, transmission frequencies, and wireless technologies. Meanwhile, a univariate approach remains beneficial for scenarios where data or computational resources are limited, enabling fast, frequency-specific forecasts that are simpler to implement for localized optimization.
  4. 4.
    We conduct experiments on frequency-selective EMF exposure datasets spanning multiple narrowband channels in the frequency range of 9 kHz–6 GHz, thus covering major network operators in the Italy and cellular network technologies. Numerical results confirm the effectiveness of the conditional over unconditional EMFusion, and multivariate over univariate EMFusion. Among various exogenous conditions, we note that the working hour condition outperforms. Furthermore, the EMFusion outperforms the best baseline by around 23.85% continuous ranked probability score (CRPS), 13.93% normalized root mean square error (NRMSE) and reduces prediction CRPS error by 22.47%.
The remainder of this paper is organized as follows. Section II presents the existing state-of-the-art for TSF in wireless networks. Section III presents the architecture of the proposed multivariate EMFusion model. The experimental setup, data analysis, and visualizations are presented in Section IV followed by numerical results and discussions in Section V. Finally, Section VI concludes the paper.
II Related Work
TSF for wireless networks has evolved from classical statistical techniques to DL and, most recently, to generative AI. This subsection provides a review of existing state-of-the-art techniques for TSF and EMF forecasting in wireless networks.
II-1 Statistical Methods for TSF
Initial approaches to EMF exposure forecasting relied on classical time-series models. These included autoregressive models like ARIMA [box2015time], as well as methods such as kernel ridge regression and support-vector regression. While computationally lightweight and interpretable, these techniques often assume stationarity and require significant manual feature engineering. Their linear parametrisation also limits their accuracy when the underlying drivers of EMF exposure interact non-linearly across multiple time scales. A representative approach is found in [de2019time], who proposed a model that decomposes wide-band EMF radiation into three distinct components: the average exposure level, daily and half-daily periodic patterns (seasonal behavior), and the temporal correlation among residuals [de2019time]. Their model effectively captured different behaviors between weekdays and weekends by applying separate autoregressive model for weekdays and Sunday, and the other autoregressive model for Saturday [de2019time]. While effective, such methods are tailored to specific observed periodicities and struggle to generalize without manual recalibration.
II-2 Deep Learning for TSF
In this subsection, we classify DL approaches for TSF in wireless networks, including network traffic forecasting, QoS prediction, channel prediction, and EMF forecasting.
Network Traffic Forecasting
[nikravesh2016mobile] compared Multi-Layer Perceptrons (MLP) as well as Support Vector Machines (SVM), to predict traffic in commercial Long-Term Evolution (LTE) networks. Similarly, [dalgkitsis2018traffic] developed a univariate forecasting platform for 4G networks using Long Short-Term Memory (LSTM) and Recurrent Neural Networks (RNNs), demonstrating superiority over statistical baselines like ARIMA. More recent studies have shifted toward multivariate and interpretable models. [di2023multivariate] characterized LTE-Advanced Voice-over-IP metrics using Vector Autoregression (VAR) and supervised DL sliding-window frameworks. To enhance interpretability, [fiandrino2024aichronolens] introduced AICHRONOLENS, an Explainable AI (XAI) framework that correlates temporal structures with LSTM error diagnoses. In the context of Software-Defined Networking, [isravel2024multivariate] utilized LSTM to model jointly evolving traffic flows for healthcare applications. Furthermore, for open radio access networks, [habib2024transformer] proposed a Transformer-based Autoformer architecture to predict aggregate cell traffic. While these methods achieve strong deterministic accuracy, they primarily focus on point forecasting. They generally lack explicit mechanisms for data imputation handling missing values and do not provide intrinsic uncertainty quantification essential for trustworthy network planning.
Channel Prediction
Channel prediction is critical for mitigating aging in high-mobility scenarios. [sone2020wireless] benchmarked various models, including Holt-Winters and XGBoost, against LSTMs for enterprise Wireless Local Area Network traffic, analyzing spatio-temporal Access Point (AP) clustering. In satellite communications, [zhang2021deep] proposed a Satellite Channel Predictor based on LSTM units for Low Earth Orbit (LEO) massive Multiple-Input Multiple-Output (mMIMO) systems. Their approach outperformed Kalman filters and standard RNNs in multi-step Channel State Information (CSI) forecasting. Similar to traffic forecasting, these works rely on deterministic models and do not account for the confidence level of predictions, which is vital for robust resource allocation.
Quality of Service (QoS) Prediction
Forecasting QoS involves modeling complex dependencies. [dinaki2021forecasting] introduced a hybrid bidirectional LSTM (BiLSTM) and Convolutional neural network (CNN) architecture for multivariate QoS forecasting in Content Delivery Networks (CDNs). [hameed2022toward] utilized a temporal Transformer encoder to predict QoS metrics in IoT networks, effectively capturing long-term dependencies that RNNs often miss. [colpitts2023short] investigated Sequence-to-Sequence (Seq2Seq) models for rural LTE Key Performance Indicators (KPIs), noting that exogenous features provided minimal gains due to short temporal dependencies.
EMF Exposure Forecasting
Research in EMF forecasting has progressed from spatial regression to time-series analysis. [kiouvrekis2024comparative] focused on comparative regression using ensemble methods like Random Forest to rank urban factors affecting EMF levels. For temporal forecasting, [pala2021examining] demonstrated that LSTMs outperform classical statistical models (e.g., ARIMA) in long-term (over a 60-month period) for univariate EMF prediction. Recently, [nguyen2024deep] demonstrated the effectiveness of 1-D CNNs in handling single-step and multi-step EMF forecasting tasks by leveraging translation invariance to identify periodicities in the data. Other research works also adapted CNN designs, utilizing architectures like Radial Basis Function Networks and Generalized Regression Neural Networks to model complex RF-EMF signals from emerging technologies like 5G [al2023machine]. Despite the aforementioned advancements, significant research gaps remain in the domain.
  1. 1.
    Existing EMF studies rely on univariate forecasting of aggregate EMF levels collected over a large bandwidth. Thus, they fail to capture the inter-operator and inter-frequency correlations essential for network planning.
  2. 2.
    Approaches like [mootoo2025emforecaster] rely on post-hoc calibration rather than learning the joint probability distribution directly. This decoupling can limit the expressiveness of the uncertainty representation. To address this, recent work has turned to distribution-free methods like conformal prediction. A notable example is EMForecaster [mootoo2025emforecaster], which integrates a powerful PatchTST backbone with a post-hoc conformal calibration layer. This two-stage pipeline delivers statistically valid prediction intervals for EMF levels. However, by decoupling uncertainty quantification from the primary sequence modeling, this approach is not directly applicable to multivariate frequency selective forecasting that adhere to the physical constraints and interdependencies across different frequency carriers, wireless technologies, and operators. Furthermore, EMForecaster [mootoo2025emforecaster] is designed for univariate forecasting, thus a large number of parallel models will be required to handle frequency selective forecasting.
  3. 3.
    Current works typically assume clean historical data, failing to address the frequent sensor outages and data gaps inherent in real-world EMF monitoring [nguyen2024deep].
II-3 Generative AI for TSF
In recent years, generative models have emerged as a powerful paradigm within DL for TSF. Representative architectures include Variational Autoencoders (VAEs) [li2022generative], Generative Adversarial Networks (GANs) [goodfellow2014generative, koochali2021if], and, more recently, DMs  [song2019generative, ho2020denoising], which have gained remarkable attention for their ability to capture data uncertainty and generate high-fidelity forecasts. DMs, as latent-variable generative frameworks, learn complex data distributions via a dual process of progressive noise injection (the forward diffusion) and iterative denoising (the reverse process). More recently, DMs have been successfully adapted to time series data [rasul2021autoregressive, tashiro2021csdi], offering a flexible generative framework for learning intricate temporal dependencies and stochastic dynamics. While DiffTCN introduces probabilistic modeling, the majority of QoS research remains deterministic.
In the context of TSF, DMs are typically formulated as conditional generative models, where the goal is to produce future trajectories conditioned on past observations and optional contextual covariates [li2022generative, rasul2021autoregressive, alcaraz2022diffusion]. During training, Gaussian noise is incrementally added to the ground-truth future sequence through the forward diffusion process, while a neural network, conditioned on historical context, is optimized to learn the inverse mapping that progressively denoises the signal. At inference, forecasts are generated by iteratively sampling and denoising noise vectors, effectively drawing from the learned conditional distribution.
Compared to conventional deterministic or probabilistic models, DMs present several advantages:
  1. 1.
    DMs yield fully probabilistic forecasts, enabling uncertainty quantification through multiple stochastic samples and predicting extreme cases in terms of different EMF traffic flows, whereas deterministic models are disadvantaged in capturing uncertainty [yan2025semantic].
  2. 2.
    They maintain stable and well-behaved training dynamics, avoiding adversarial divergence;
  3. 3.
    They offer a highly modular design that allows flexible conditioning on exogenous factors such as seasonal indicators, control variables, or multimodal inputs [kollovieh2023predict, yuan2024diffusion, liu2024retrieval].
Empirically, DMs demonstrate competitive or even state-of-the-art results across numerous TSF benchmarks [shen2024multi, liu2024retrieval, li2025diffusion, meijer2024rise, su2025diffusion], positioning them as a cornerstone of next-generation generative forecasting frameworks.
Figure 1: The U-Net architecture used for EMFusion. The network processes a noisy input through a symmetric encoder-decoder path. Timestep embeddings are incorporated into the residual blocks, while external context is injected via a cross-attention block between the residual blocks of the encoder and decoder. The residual block architecture of EMFusion to estimate $\epsilon_{\theta}(\cdot)$.
III Multivariate EMFusion: Conditional DM for Frequency Selective EMF Forecasting
This section details our methodology for multivariate EMF TSF leveraging a conditional DM. Given the sequence of historical EMF measurements and corresponding contextual factors, the goal is to model the conditional distribution of the future EMF values. The proposed model is then enhanced with imputation-based resampling to deal with real-world EMF logs where EMF measurements may be irregularly sampled or there are gaps or sensor outages.
III-A Multivariate Conditional Diffusion Model
Figure 2: EMFusion for multivariate frequency selective EMF forecasting with cross-attention.
We consider a multivariate time-series forecasting problem in which each observation $\mathbf{x}_{t}\in\mathbb{R}^{N}$ represents measurements from $N$ variates or frequency channels at time $t$. Given a historical window of length $H$, denoted by $\mathbf{X}_{\text{past}}=(\mathbf{x}_{T-H+1},\ldots,\mathbf{x}_{T})$, the objective is to predict the future multichannel trajectory $\mathbf{X}_{\text{future}}=(\mathbf{x}_{T+1},\ldots,\mathbf{x}_{T+F})$ over a forecast horizon $F$. To model the conditional distribution $p_{\theta}(\mathbf{X}_{\text{future}}\mid\mathbf{X}_{\text{past}},\mathbf{c})$, we introduce an external condition vector $\mathbf{c}$, which may include calendar features, operational metadata, or other exogenous covariates. Unlike vectorized formulations, we treat the full future window as a matrix $\mathbf{X}_{0}:=[\mathbf{X}_{\text{past}},\mathbf{X}_{\text{future}}]\in\mathbb{R}^{(H+F)\times N}$ and apply diffusion directly in this multivariate tensor space. This enables probabilistic forecasting while preserving the temporal–channel structure of the signal.
In the diffusion-based formulation, the forward noising process corrupts the clean $\mathbf{X}_{0}$ over $T$ time steps according to the Markov transition process given by [ho2020ddpm, huang2025probabilistic]:
$q(\mathbf{X}_{t}\mid\mathbf{X}_{t-1})=\mathcal{N}\!\left(\sqrt{\alpha_{t}}\,\mathbf{X}_{t-1},(1-\alpha_{t})\mathbf{I}\right),$ (1)
which admits the closed-form marginal
$q(\mathbf{X}_{t}\mid\mathbf{X}_{0})=\mathcal{N}\!\left(\sqrt{\bar{\alpha}_{t}}\,\mathbf{X}_{0},\,(1-\bar{\alpha}_{t})\mathbf{I}\right),$ (2)
where $\alpha_{t}\in(0,1]$ is a predefined noise schedule, $\mathbf{I}$ denotes the identity matrix, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ , where $\alpha_{s}$ is the noise intensity of each step in the noise addition process. During training, noisy future samples can be directly sampled from $\mathbf{X}_{0}$ at any timestep $t$, i.e.,
$\mathbf{X}_{t}=\sqrt{\bar{\alpha}_{t}}\,\mathbf{X}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon},\qquad\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),$ (3)
with $\bm{\epsilon}\in\mathbb{R}^{(H+F)\times N}$ containing i.i.d. Gaussian noise.
The training objective minimizes the discrepancy between the true noise $\bm{\epsilon}$ and the network-predicted noise $\bm{\epsilon}_{\theta}(\mathbf{X}_{t},t,\mathbf{c})$ using a mean-squared error loss:
$L(\theta)=\mathbb{E}_{t,\mathbf{X}_{0},\bm{\epsilon}}\!\left[\big\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\!\left(\sqrt{\bar{\alpha}_{t}}\,\mathbf{X}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon},t,\mathbf{c}\right)\big\|^{2}\right].$ (4)
The reverse diffusion process serves as the generative mechanism for forecasting. The conditional reverse diffusion process predicts the distribution $p_{\theta}(\mathbf{X}_{t-1}\mid\mathbf{X}_{t},\mathbf{c})$, of the less noisy state $\mathbf{X}_{t-1}$ given the current noisy state $\mathbf{X}_{t}$ and the conditioning information $\bm{c}$. It learns to reconstruct the original signal $\mathbf{X}_{0}$ by iteratively denoising the noisy samples, starting from pure noise $\mathbf{X}_{T}\sim\mathcal{N}(0,\mathbf{I})$. A neural network $\bm{\epsilon}_{\theta}(\mathbf{X}_{t},t,\bm{c})$ is employed to approximate the reverse transition distribution $p_{\theta}(\mathbf{X}_{t-1}|\mathbf{X}_{t},\mathbf{c})$ by predicting the noise component $\bm{\epsilon}$ that was added at step $t$. This noise prediction network takes the noisy data $\mathbf{X}_{t}$, the current timestep $t$, and the conditioning vector $\mathbf{c}$ to predict the following conditional distribution:
$p_{\theta}(\mathbf{X}_{t-1}\mid\mathbf{X}_{t},\mathbf{c})=\mathcal{N}\!\bigl(\bm{\mu}_{\theta}(\mathbf{X}_{t},t,\mathbf{c}),\bm{\Sigma}_{t}(\mathbf{X}_{t},t,\mathbf{c})\bigr),$ (5)
where the mean computed from $\bm{\epsilon}_{\theta}$ is given by:
$\bm{\mu}_{\theta}(\mathbf{X}_{t},t,\mathbf{c})=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{X}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\bm{\epsilon}_{\theta}(\mathbf{X}_{t},t,\mathbf{c})\right).$ (6)
and $\mathbf{c}$ encapsulates features from the dataset, such as Italy_WorkingDay, Italy_WorkingHour, and Italy_Season, which depict factors influencing EMF variations [meijer2024rise, su2025diffusion]. By conditioning the noise prediction on $\mathbf{c}$, the model learns to generate EMF trajectories consistent with temporal continuity and relevant external dependencies.
This formulation enables a fully multivariate, conditionally guided reverse diffusion process in which $\mathbf{X}_{t}$ is iteratively denoised using both the historical multichannel window $\mathbf{X}_{\text{past}}$ and the external condition vector $\mathbf{c}$.
During inference, a classifier-free guidance (CFG) mechanism [ref:classifier-free, huang2025probabilistic] steers the generation process. The final noise prediction is obtained by combining the conditional and unconditional outputs of the denoising network:
$\hat{\bm{\epsilon}}_{\theta}(\mathbf{X}_{t},t,\mathbf{c})=(1+s_{c})\,\bm{\epsilon}_{\theta}(\mathbf{X}_{t},t,\mathbf{c})-s_{c}\,\bm{\epsilon}_{\theta}(\mathbf{X}_{t},t,\emptyset),$ (7)
where $s_{c}$ is the guidance scale controlling the adherence to the conditioning input $\bm{c}$, and $\emptyset$ represents an unconditional prediction. In conditional TSF, positive CFG scales often over-concentrate the posterior, which narrows prediction intervals. Since our model already injects conditioning via cross-attention, we found $s_{c}=0$ is the best via hyperparameter tuning. This conditional diffusion formulation provides the theoretical basis for our model.
Input: $\mathbf{X}_{0}=[\mathbf{X}_{\text{past}},\mathbf{X}_{\text{future}}]$, conditions $\mathbf{c}$
Output: Learned parameters $\theta$
Sample $t\sim\mathcal{U}\{1,\ldots,T\}$ and noise $\bm{\epsilon}\sim\mathcal{N}(0,I)$;
Generate noisy sequence: $\mathbf{X}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{X}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$;
Pass $\mathbf{X}_{t}$ and $t$ through U-Net;
Time series $t$ embedding
for each cross-attention block do
  Positional encoding computed by Eq (15);
   Compute $Q=\mathbf{h}_{\mathrm{flat}}W_{Q}$, $K=\mathbf{c}W_{K}$, $V=\mathbf{c}W_{V}$;
   Compute $\text{Attention}=\mathrm{Softmax}(QK^{\top}/\sqrt{e_{k}})V$;
   Update feature map: $\mathbf{h}\leftarrow\mathbf{h}+\text{Attn}$;
end for
Predict noise with Cross Attention by Eq (12): $\hat{\bm{\epsilon}}_{\theta}=\epsilon_{\theta}(\mathbf{X}_{t},t,\mathbf{c})$;
Update parameters by minimizing $\mathcal{L}(\theta)=\|\bm{\epsilon}-\hat{\bm{\epsilon}}_{\theta}\|_{2}^{2}$;
Algorithm 1 DDPM Training for Multivariate Forecasting with Cross-Attention
Input: $\mathbf{X}_{\mathrm{past}}$, mask $\bm{\Omega}$, conditions $\mathbf{c}$, unknown future window $\mathbf{X}_{\mathrm{future}}$
Output: Sampled future trajectory $\widehat{\mathbf{X}}_{\mathrm{future}}$
Initialize $\mathbf{X}_{T}\in\mathbb{R}^{(H+F)\times N}\sim\mathcal{N}(0,I)$, $\mathbf{X}_{0}=[\mathbf{X}_{\mathrm{past}},\mathbf{X}_{\mathrm{future}}]$;
for $t=T$ to 11 do
   Time series $t$ embedding
   Positional encoding computed by Eq (15);
   Predict noise with Cross Attention by Eq (12); $\hat{\bm{\epsilon}}_{\theta}=\epsilon_{\theta}(\mathbf{X}_{t},t,\mathbf{c})$;
   Sample $\mathbf{z}\sim\mathcal{N}(0,I)$ if $t>1$, else $\mathbf{z}=0$;
   Compute one step denoised estimate: $\mathbf{X}_{t-1}^{\prime}=\frac{1}{\sqrt{\alpha_{t}}}\!\left(\mathbf{X}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\hat{\bm{\epsilon}}_{\theta}\right)+\sigma_{t}\mathbf{z}$;
   Sample $\mathbf{\epsilon}\sim\mathcal{N}(0,I)$ for $\mathbf{X}_{\mathrm{past}}$ if $t>1$, else $\mathbf{\epsilon}=0$;
   Add noise to $\mathbf{X}_{\mathrm{0}}$: $\mathbf{X}_{\text{obs}}=\sqrt{\bar{\alpha}_{t}}\,\mathbf{X}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon}$;
   Clamp observed values (imputation step): $\mathbf{\hat{X}}_{t-1}=\bm{\Omega}\odot\mathbf{X}_{\mathrm{obs}}+(1-\bm{\Omega})\odot\mathbf{X}_{t-1}^{\prime}$;
end for
Extract $\widehat{\mathbf{X}}_{\mathrm{future}}$ from $\mathbf{X}_{0}$;
Algorithm 2 DDPM Inference for Multivariate Forecasting with Cross-Attention and Masked Imputation
TABLE II: Frequency $f$ (MHz) allocation for mobile operators and technology band used in the Italian dataset. Each entry shows the antenna factor $(\alpha_{f})$ in dB m-1 from Keysight N6850A calibration [keysight_n6850a].
Operators $\mathcal{F}_{\mathrm{Iliad}}$ 763.0 (32.8), 1835.0 (31.2), 2150.0 (31.9), 2635.0 (33.2), 3630.0 (36.6)
$\mathcal{F}_{\mathrm{TIM}}$ 773.0 (32.9), 806.0 (33.3), 935.0 (34.7), 2137.0 (31.8), 2137.5 (31.8), 2662.0 (33.3), 2662.5 (33.3), 3760.0 (36.9), 3468.0 (36.0), 3468.5 (36.0), 3469.0 (36.0), 3568.0 (36.3), 3568.5 (36.3)
$\mathcal{F}_{\mathrm{VF}}$ 783.0 (33.0), 816.0 (33.4), 945.0 (34.8), 1870.0 (31.0), 2162.0 (31.9), 2162.5 (31.9), 2647.0 (33.2), 2647.5 (33.2), 3680.0 (36.7)
$\mathcal{F}_{\mathrm{W3}}$ 796.0 (33.2), 955.0 (34.9), 1850.0 (30.9), 2120.0 (31.7), 2585.0 (33.0), 2680.0 (33.4), 3610.0 (36.5)
Technologies $\mathcal{F}_{\mathrm{2G}}$ 935.0 (34.7), 945.0 (34.8), 955.0 (34.9)
$\mathcal{F}_{\mathrm{3G}}$ 2120.0 (31.7), 2137.0 (31.8), 2137.5 (31.8), 2150.0 (31.9), 2162.0 (31.9), 2162.5 (31.9)
$\mathcal{F}_{\mathrm{4G}}$ 763.0 (32.8), 773.0 (32.9), 783.0 (33.0), 796.0 (33.2), 806.0 (33.3), 816.0 (33.4), 1835.0 (31.2), 1850.0 (30.9), 1870.0 (31.0), 2585.0 (33.0), 2635.0 (33.2), 2647.0 (33.2), 2647.5 (33.2), 2662.0 (33.3), 2662.5 (33.3), 2680.0 (33.4)
$\mathcal{F}_{\mathrm{5G}}$ 3468.0 (36.0), 3468.5 (36.0), 3469.0 (36.0), 3568.0 (36.3), 3568.5 (36.3), 3610.0 (36.5), 3630.0 (36.6), 3680.0 (36.7), 3760.0 (36.9)
III-B Model Architecture and Cross-Attention Integration
We employ a U-Net backbone for the noise-prediction network $\bm{\epsilon}_{\theta}$. As shown in Fig. 1, the U-Net is a deep neural network with multiple residual blocks. The encoder and decoder are symmetric structure, consisting of the same number of residual blocks and additional downsampling or upsampling block. Each residual block contains convolution to capture local details of the long series, with the input $\mathbf{X_{t}}$, condition $\mathbf{c}$, and time step $t$. The input multivariate $\mathbf{X}_{0}$ passes through the U-Net and is processed by residual block to an intermediate feature map $\mathbf{h}$ (with dimensions $C_{\mathbf{h}}\times h_{\mathbf{h}}\times w_{\mathbf{h}}$), while $\mathbf{h}=\mathbf{X}_{t}$ in the first layer. Here, $C_{\mathbf{h}}$ is the number of channels in the convolutional layer, $h_{\mathbf{h}}$ and $w_{\mathbf{h}}$ are the spatial height and width of the feature map, respectively.
III-B1 Residual Block with Time-Step Embedding
The core of the $\epsilon_{\theta}$ network is a stack of residual blocks that progressively refine the feature maps. As shown in Fig. 1, each residual block applies normalization, nonlinearity, convolution, and time step conditioning, followed by a residual shortcut. Given an input feature map $\mathbf{h}$ from the previous layer, the block first applies Group Normalization and a Sigmoid Linear Unit (SiLU) activation, followed by a $3\times 3$ convolution, i.e., [ronneberger2015u]:
$\tilde{\mathbf{h}}=\mathrm{Conv}_{1}\!\bigl(\sigma(\mathrm{GN}(\mathbf{h}))\bigr),$ (8)
where $\mathrm{GN}(\cdot)$ denotes GroupNorm, $\sigma(\cdot)$ is the SiLU activation, and $\mathrm{Conv}_{1}$ is a $3\times 3$ convolution.
A critical component of the block is the injection of the diffusion step embedding. The global timestep embedding $\mathbf{t}^{e}$ is passed through a block-specific linear layer and then added to the intermediate feature map:
$\mathbf{h}_{i}=\tilde{\mathbf{h}}+\mathrm{Linear}_{t}\!\bigl(\sigma({{\mathbf{t}^{e}}})\bigr),$ (9)
where $\mathrm{Linear}_{t}$ maps ${{\mathbf{t}^{e}}}\in\mathbb{R}^{d_{\text{embed}}}$ to the channel dimension of the block, and the result is broadcast across spatial dimensions. This makes the block’s behavior explicitly adaptive to the current diffusion step $t$.
To incorporate the diffusion timestep to the residual block and enrich its effective information, a sinusoidal time embedding is employed to expand its dimension. The timestep $t$ is first mapped into a positional embedding vector $\mathbf{t}^{e}$ as
$\mathbf{t}^{e}=\text{Concat}\Big($ $\left[\cos(t\cdot\omega_{0}),\dots,\cos\!\left(t\cdot\omega_{\frac{d}{2}-1}\right)\right],$ (10)
$\left[\sin(t\cdot\omega_{0}),\dots,\sin\!\left(t\cdot\omega_{\frac{d}{2}-1}\right)\right]\Big)$
For a given dimension index $m$ where $0\leq m<C_{\mathbf{h}}/2$, the frequencies are defined as follows:
$\omega_{m}=\exp(-\ln(10000)\cdot\frac{2m}{C_{\mathbf{h}}})$ (11)
We integrate condition $\mathbf{c}$ into the feature map using cross-attention operation as shown below (more details of the cross-attention are given in the following subsection and Fig. 2):
$\hat{\mathbf{h}}=\mathbf{h}_{i}+\text{CrossAttention}(\mathbf{h},\mathbf{c}).$ (12)
The updated feature map $\hat{\mathbf{h}}$ is then processed by a second sequence of operations, GroupNorm, SiLU, dropout, and another $3\times 3$ convolution [ronneberger2015u]:
[…]
Finally, a residual shortcut is applied as [he2016deep]:
$\mathbf{h}_{\text{out}}=\mathbf{h}^{\prime}+\mathrm{Shortcut}(\mathbf{h}),$ (14)
where $\mathrm{Shortcut}$ is the identity mapping for feature mining and convolution operation for up sampling and down sampling. By stacking these modules in the encoder, the series is processed shorter and the convolution channels increase, so that abstract temporal and spatial information are extracted.
III-B2 Cross-Attention Module
To integrate conditioning information, instead of simple concatenation, we use cross-attention modules at multiple layers in both the encoder and decoder, as depicted in Fig. 2. This allows the model to dynamically attend to the most relevant parts of the conditioning information at different stages of processing and across different temporal scales. The condition tensor $\mathbf{c}$ (with dimensions $N_{f}\times d_{f}$) and the intermediate feature map $\mathbf{h}$ are two inputs of the cross-attention module. Note that $N_{f}$ is the sequence length of the condition and $d_{f}$ is its feature dimension.
To compute the attention, the 3D feature map $\mathbf{h}$ is first flattened into a 2D tensor, $\mathbf{h}_{\text{flat}}$, with shape $(C_{h},N_{h})$, where $N_{h}=h_{h}\cdot w_{h}$. Since cross-attention calculation basically contains linear connection without recurrence, it’s important to fully make use of the order of the sequence. Positional encoding adds positional information to $\mathbf{c}$ and $\mathbf{h}_{\text{flat}}$, which helps the module extract temporal and spatial correlations. Consequently, we add positional encoding to $\mathbf{c}$ and $\mathbf{h}_{\text{flat}}$ according to [ref:attention]:
$\mathbf{c}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\leftarrow}\mathbf{c}+PE(\mathbf{c})\in\mathbb{R}^{N_{f}\times d_{f}}$ (15)
$\mathbf{h}_{\text{flat}}\leftarrow\mathbf{h}_{\text{flat}}+PE(\mathbf{h}_{\text{flat}})\in\mathbb{R}^{C_{h}\times N_{h}}$ (16)
$PE(\cdot)=\begin{cases}PE_{(p,2i)}=sin(p/10000^{2i/d_{\mathrm{series}}}),\\ PE_{(p,2i+1)}=cos(p/10000^{2i/d_{\mathrm{series}}}),\end{cases}$ (17)
where $p$ and $i$ denote the position and the dimension index of the input series, respectively, and $d_{\text{series}}$ is the embedding dimension of the sequence.
The cross-attention module operates between the feature map and the conditional embeddings, allowing the model to explicitly learn their correlations and dynamically modulate its representations based on the provided conditions. Its calculation is based on $QKV$ where the Queries ($Q$) are projected from this flattened feature map, while the Keys ($K$) and Values ($V$) are projected from the conditioning tensor $\mathbf{c}$, i.e.,
$Q=\mathbf{h}_{\text{flat}}W_{Q},\quad K=\mathbf{c}W_{K},\quad V=\mathbf{c}W_{V}$ (18)
where $W_{Q}$, $W_{K}$, and $W_{V}$ are learnable weight matrices [ref:latent, huang2025probabilistic]. The attention matrix $A$ and the final output are calculated using the standard scaled dot-product attention mechanism:
$\text{CrossAttention}(\mathbf{h},\mathbf{c})=\text{Softmax}\!\left(\frac{QK^{T}}{\sqrt{e_{k}}}\right)V=AV,$ (19)
where $e_{k}$ is the dimension of the keys (and queries), used for scaling. According to ① in Fig. 2, a single row of the attention matrix $A$ represents the set of attention weights that one specific element from the flattened feature map $\mathbf{h}_{\text{flat}}$ (represented by a row of $Q$) pays to all $N_{f}$ elements of the condition vector $\mathbf{c}$ (represented by the rows of $K$). ② in Fig. 2 shows a column of the value $V$ that corresponds to a feature map weighted by the attention in $A$ [petit2021u]. The output of this operation, $AV$, represents the contextually relevant information from $\mathbf{c}$, weighted according to its relevance to $\mathbf{h}$. This output is then reshaped and projected to match the original feature map’s dimensions ($C_{h}\times h_{h}\times w_{h}$). This result is finally added back to the original feature map $\mathbf{h}$ via a residual connection, which is shown in the previous section:
$\mathbf{h}^{\prime}=\mathbf{h}+\text{CrossAttention}(\mathbf{h},\mathbf{c}).$
In this context, $\text{CrossAttention}(\cdot)$ in Eq. (12) represents the full module’s operation, including the final reshape and projection steps shown in the diagram. According to ② in Fig. 2, a single column of the final output ($AV$) is the new vector representation for that specific feature map element. It is calculated as the weighted sum of all $N_{f}$ rows of the Value matrix $V$. The weights used for this sum are the attention scores from the corresponding row of $A$.
III-C EMFusion Training and Inference
The cross-attention integrated U-Net learns to predict noise in the training phase as Algorithm 1. Given the full window $\mathbf{X}_{0}=[\mathbf{X}_{\text{past}},\mathbf{X}_{\text{future}}]$ and the conditions $\mathbf{c}$, the first step is to randomly choose a $t\sim\mathcal{U}\{1,\ldots,T\}$ and noise $\bm{\epsilon}\sim\mathcal{N}(0,I)$ to add noise to the original sequence $\mathbf{X}_{0}$ by using Eq. (3): $\mathbf{X}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{X}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$. The task of the U-Net is to learn the ability to predict $\mathbf{\epsilon}$ given $\mathbf{X}_{t},t,\mathbf{c}$. Pass $\mathbf{X}_{t}$ through U-Net and generate intermediate feature map $\mathbf{h}$. Then embed positional information to $\mathbf{c}$ and $\mathbf{X}_{0}$ and calculate cross-attention between $\mathbf{h}$ and $\mathbf{c}$ when going through cross-attention blocks. Subsequently the output noise $\hat{\bm{\epsilon}}_{\theta}$ is the same dimension as $\mathbf{\epsilon}$. By minimizing MSE loss $\mathcal{L}(\theta)=\|\bm{\epsilon}-\hat{\bm{\epsilon}}_{\theta}\|_{2}^{2}$, the model updates itself and learn to predict noise based on any $t$. Then the well-trained model participates in the denoising process in the inference phase by predicting noise $\hat{\bm{\epsilon}}_{\theta}=\epsilon_{\theta}(\mathbf{X}_{t},t,\mathbf{c})$ at each step $t$ and the conditions are input through cross-attention module, as shown in Algorithm 2.
III-D Imputation-Based Sampling at the Inference
In practical EMF monitoring, historical measurements are often uneven, partially observed, or corrupted due to sensor outages, data-handling issues, and privacy-related filtering, resulting in time series with substantial gaps and irregular sampling. Such incomplete inputs limit the effectiveness of pure conditional forecasting, which assumes fully reliable past observations. An imputation-based diffusion formulation addresses this limitation by treating both missing past values and all future values as unobserved entries of a single masked sequence. During the DDPM reverse process, the model performs principled generative impainting—reconstructing irregular or missing segments of the past while simultaneously generating future values under the same probabilistic dynamics. Since past and future EMF measurements lie on the same underlying temporal manifold, casting forecasting as structured inpainting enables the model to learn cross-temporal dependencies across the entire sequence.
To handle both forecasting and imputation within a unified framework, we adopt a masked conditional DDPM that models the full sequence $\mathbf{X}_{\text{0}}=[\mathbf{X}_{\text{past}},\mathbf{X}_{\text{future}}]$ under partial observations. That is, the model receives a concatenated input sequence $\mathbf{X}_{\text{0}}$ consisting of a known historical window $\mathbf{X}_{\text{past}}$ and an unknown future window $\mathbf{X}_{\text{future}}$ that is initially masked. The objective is to infer the missing future segment in a manner that is probabilistically consistent with $\mathbf{X}_{\text{past}}$. The algorithmic realization is summarized in Algorithm 2.
Let $\bm{\Omega}$ denotes a binary observation mask (for the i-th element, $\bm{\Omega}_{i}=1$ for observed entries and $\bm{\Omega}_{i}=0$ for missing ones), and let $\mathbf{X}_{\text{obs}}=\sqrt{\bar{\alpha}_{t}}\,\mathbf{X}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),$ denotes the available measurements. As in standard diffusion models, the forward process gradually perturbs the clean sequence $\mathbf{X}_{0}$ with Gaussian noise as in (3), while the model learns a masked conditional reverse process, i.e., (5) still applies by adding conditions on $\mathbf{c}$. The noisy sequence is processed through the conditional U-Net $\bm{\epsilon}_{\theta}(\mathbf{X}_{t},t,\mathbf{c})$ to obtain a one-step denoised estimate of the entire trajectory, denoted as $\hat{\mathbf{X}}_{t-1}$. After each denoising step, the observed entries are clamped to their known values to ensure consistency:
$\mathbf{\hat{X}}_{t-1}=\mathbf{\bm{\Omega}}\odot\mathbf{X}_{\text{obs}}+(1-\mathbf{\bm{\Omega}})\odot\mathbf{{X}}^{\prime}_{t-1},$ (20)
where $\hat{X}_{t-1}$ denotes the model’s predicted denoised sample at timestep t1t-1.
$\hat{\mathbf{X}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\!\left(\mathbf{X}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\hat{\bm{\epsilon}}_{\theta}\right)+\sigma_{t}\mathbf{z}$ (21)
After $T$ iterations of the denoising process, a predicted sequence $\widehat{\mathbf{X}}_{\mathrm{future}}$ can be obtained from $\mathbf{\hat{X}}_{0}$. This mechanism enables the DM to impute with past values and generate future trajectories within the same sampling procedure. The training objective follows the standard noise-prediction loss as in (4) with masked conditioning. If we sample $\hat{N}$ times, a set of $\hat{N}$ predicted sequences $\mathbf{S}=\{\hat{\mathbf{X}}_{\mathrm{future}}^{i}\}_{i=1,\cdots,\hat{N}}$ is ready for probabilistic interval construction which is explained next.
III-E Probabilistic Interval Construction
Subsection III-E describes probabilistic interval construction for calibrated uncertainty quantification. The procedure begins once the EMFusion has generated an ensemble of $\hat{N}$ distinct future EMF trajectories $\mathbf{S}=\{\hat{\mathbf{X}}_{\mathrm{future}}^{i}\}_{i=1,\cdots,\hat{N}}\in\mathbb{R}^{\hat{N}\times F}$ for each of the $N$ monitored carriers (e.g., individual 2G, 3G, 4G, and 5G carriers across multiple operators). For each future time step $k\in\{1,2,\cdots,F\}$, this yields a set of $\hat{N}$ predicted values, denoted as $\mathbf{Z}=\{z_{1},z_{2},\dots,z_{\hat{N}}\}$. This ensemble can be viewed as a discrete, non-parametric sample drawn from the learned conditional predictive distribution. To convert this sample set into a continuous probability density function (PDF), Kernel Density Estimation (KDE) is employed [huang2025probabilistic]. KDE is a classical non-parametric technique that estimates the underlying density $f(z)$ by placing a kernel function, typically Gaussian, at each sample point and summing the contributions:
$f(z)=\dfrac{1}{\hat{N}\cdot s_{p}}\sum\limits_{i=1}^{\hat{N}}K\left(\dfrac{z-z_{i}}{s_{p}}\right),$ (22)
where $K(\cdot)$ is the kernel function and $s_{p}$ is the bandwidth parameter. The bandwidth $s_{p}$ is crucial for controlling smoothness, determining the bias and variance trade-off in the estimated density. This nonparametric statistical method can directly estimate probability distributions from discrete scenarios without requiring prior assumptions about the distribution. Once the PDF $f(z)$ is derived, the cumulative distribution function $F(z)$ is computed, and the prediction interval is obtained via quantile extraction. Specifically, for a target confidence level $\gamma$, the lower and upper quantile bounds $q_{\underline{\alpha}}$ and $q_{\overline{\alpha}}$ satisfy $PI_{\gamma}=[q_{\underline{\alpha}},q_{\overline{\alpha}}]$, $\qquad q_{\alpha}=F^{-1}(\alpha)$, where $\gamma=\overline{\alpha}-\underline{\alpha}$, and $\underline{\alpha}$, $\overline{\alpha}$ denote the lower and upper quantile levels, respectively.
In most KDE-based interval estimation scenarios, symmetric intervals satisfying $\overline{\alpha}+\underline{\alpha}=1$ are adopted. In this work, however, since EMFusion directly generate multiple stochastic trajectories, we employ a simplified empirical interval construction approach. Specifically, given sorted samples $z_{1}<z_{2}<\dots<z_{\hat{N}}$, the interval $PI_{\gamma}$ is directly determined from the corresponding order statistics without KDE, offering a computationally efficient alternative:
$\hat{N}_{\gamma}=\lfloor(1-\gamma)\cdot\hat{N}\rfloor$ (23)
$PI_{\gamma}=[z_{{\hat{N}}_{\gamma}},z_{\hat{N}-\hat{N}_{\gamma}}]$ (24)
III-F Univariate Forecasting: A Special Case
Figure 3: Comparison of univariate (UV) and multivariate (MV) forecasting strategies for EMF time-series across NN features.
As illustrated in Fig. 3, Univariate forecasting illustrates the aggregate-then-learn paradigm, where multiple time series are first aggregated before applying the forecasting model. Multivariate forecasting the learn-then-aggregate paradigm, where forecasts are generated for each frequency series individually and then aggregated. In the EMFusion framework, the main differences between the univariate and multivariate instantiations of EMFusion are confined to the input dimensionality and the associated architectural hyperparameters. In the multivariate setting ($N>1$), the input $\mathbf{X}_{0}$ is treated as a 2-D tensor, processed via 2-D convolutional layers to simultaneously capture temporal evolution and inter-frequency correlations. Conversely, the univariate configuration ($N=1$) utilizes 1-D convolutions, focusing exclusively on temporal dynamics. In the univariate setting, the model receives a single narrow-band EMF time series and the input tensor is of shape $(B,H,1)$, where $B$ is the batch size and $H$ is the look-back length, thus we employ 1-D convolutions. In the multivariate setting, the input at each time step lead to an input tensor of shape $(B,H,N)$ and 2-D convolutions acting jointly on time and frequency. When computing cross-attention, the univariate uses a single hidden channel $h_{h}=1$ in the value projection (effectively attending only along time), whereas the multivariate sets $h_{h}>1$ so that attention can propagate information both temporally and across bands.
IV Data Specifications and Analysis
This section details dataset specifications including the measurement set-up, received power (dBm) to electric-field (Volts per meter) conversion, and the external conditioning factors.
Figure 4: EMF levels in Italian dataset, as a function of network operators (1,2) and cellular technologies (3,4). The plots (1,3) display the 8-month trend (Aug 2023 – Mar 2024), while the plots (2,4) show enlarged 14-day daily trend.
The top two plots in Fig. 4 show the time evolution [^1] The data were collected at the University Hospital Tor Vergata in Rome, Italy, from 2023-07-26 16:49:02 to 2024-03-11 05:08:28. of operator–specific EMF levels in the Italian dataset. The operator-wise difference in EMF levels is notable. Iliad systematically dominates the total exposure, with a typical peak at $0.8$ v/m. W3 provides the second-largest contribution on EMF, while TIM and VF remain consistently lower, with most values below $0.3$ V/m. This suggests antenna being placed in the University hospital or a configuration that produced a stronger signal at the specific point of measurement in favour of Iliad. Moreover, while the other operators display broadly similar deployment and load patterns, Iliad follows a more independent deployment and load profile, with a narrower spectrum allocation in the mid-band 5G frequencies compared to the other operators.
On the other hand, the bottom two plots in Fig. 4 depict that the EMF level from 4G clearly dominates the total EMF levels from 2G, 3G, or 5G technologies, both in terms of average level and short–term variability, with typical peaks reaching $0.8$ V/m. The 3G contribution is noticeably smaller but still non-negligible, while 2G and 5G produce the lowest fields, generally below $0.1$ V/m except for a few isolated peaks. This trend is likely due to the sparse deployment of both 2G and 5G technologies and widespread deployment of 4G at the time of measurements. Both operator-wise and technology-wise display the recurrent daily cycle and stable behaviour over the months. According to Fig. 5, 4G is only weakly correlated with the other technologies ($\rho\leq 0.19$), whereas 5G exhibits a moderate correlation with 2G and 3G. This behavior is consistent with the technological innovations introduced in 4G and 5G with respect to legacy generations, such as active antennas, beamforming, MIMO, and time-division duplexing.
Figure 5: Correlation maps of EMF levels for (a) frequency bands, (b) cellular technologies, and (c) network operators.
IV-A Dataset Specifications
The received power measurements were collected continuously from July 26, 2023 to March 10, 2024 across 29 frequency channels spanning from 9 kHz - 6GHz. The dataset is then organized according to the predefined frequency bands of major Italian network operators and the corresponding cellular technologies. Table II summarizes the frequency allocations used by each operator and technology.
IV-A1 EMF Monitoring Set-up
The EMF monitoring chain couples an omnidirectional antenna, an Anritsu MS27102A remote spectrum-monitoring unit, and the ARPA-Lazio Search24 control software [chiaraviglio2022six]. The Keysight N6850A passive antenna [keysight_n6850a] is connected to the portable spectrum analyzer via a 1.5 m low-loss coaxial cable. The antenna delivers uniform gain across all licensed cellular bands up to the 3.6 GHz. The weather-proof MS27102A sweeps 9 kHz–6 GHz continuously and, thanks to its internal pre-selector, outputs real-time narrow-band channel power while operating unattended on a rooftop site [ms27102a]. Search24 [search24] scripts enable streaming the data to a local computer over a dedicated Ethernet link.
IV-A2 Conversion of Received Power to EMF
The received power for a given frequency channel $f$ at the output of the antenna is measured in dBm. For a given frequency $f$, the received power is then converted into the electric-field strength (measured in V/m) as given below [chiaraviglio2021health]:
$E_{f}=\sqrt{\frac{P_{\mathrm{r}}(f)\,Z_{0}}{A_{\mathrm{e}}(f)}}=\sqrt{\frac{10^{\frac{P_{\mathrm{dBm}}(f){-}30}{10}}Z_{0}}{A_{\mathrm{e}}(f)}}$ (25)
where $Z_{0}\!\approx\!376.73~\Omega$, which is the intrinsic (wave) impedance of free space. For a measurement antenna with linear gain $G_{f}$, the effective aperture $A_{\mathrm{e}}(f)$ is defined as follows:
$A_{\mathrm{e}}(f)=\frac{G_{f}\lambda_{f}^{2}}{4\pi},\quad\lambda_{f}=\frac{c}{f},\quad G_{f}=\left(\frac{A_{GC}}{\lambda_{f}\cdot 10^{\alpha_{f}/20}}\right)^{2}$ (26)
where $c$ is the speed of light. $A_{GC}$ is antenna-gain calibration constant. The antenna factor $\alpha_{f}$ changes as a function of frequency $f$ as in [chiaraviglio2022six, keysight_n6850a_antfac].
IV-A3 External Conditions
User activity patterns, driven by work schedules and lifestyle habits, directly influence device usage and consequently the received power and EMF levels. To account for these temporal variations, calendar-based contextual information is incorporated as an external conditioning signal for EMF prediction. For each observation with time stamp $t$, three features were appended:
  • Italy_WorkingDay $\in\{0,1\}$: indicates whether $t$ falls on a weekday (Monday–Friday) and is not listed in the Italian public–holiday calendar. [^2] The holiday list covers national and widely observed regional holidays for 2023–2024, including Easter, Liberation Day, Labour Day, Republic Day, Ferragosto, All Saints’ Day, Immaculate Conception, and the Christmas–New Year period. It also includes the university closure date and no teaching activities dates for Università degli Studi di Roma Tor Vergata.[TorVergataAcademicCalendarPage2025]
  • Italy_WorkingHour $\in\{0,1\}$: equals 1 when the local clock time lies inside the standard business window [09:00,17:00)[09{:}00,17{:}$0$) and 0 otherwise.
  • Italy_Season $\in\{1,2,3,4\}$: a flag assigned based on the astronomical equinox–solstice boundaries, i.e., 1) Spring: March 21 to June 20; 2) Summer: June 21 to September 22; 3) Autumn: September 23 to December 20; and 4) Winter: December 21 to March 20 in the next year.
  • Italy_Multi $\in\{1,2,3,4\}\times\{0,1\}$: a two-dimensional multi-condition feature obtained by concatenating Italy_Season and Italy_WorkingDay.
IV-B EMF Classification and Correlation Analysis
From Table II, the aggregate EMF associated with any given operator or cellular technology can be obtained as follows:
$E_{\mathcal{F}_{\chi}}(t)=\biggl(\sum_{f\in\mathcal{F}_{\chi}}E^{2}_{f}\biggr)^{1/2},$ (27)
where $\mathcal{F}_{\chi}\in\{\mathrm{2G},\mathrm{3G},\mathrm{4G},\mathrm{5G},\mathrm{TIM},\mathrm{VF},\mathrm{W3},\mathrm{Iliad}\}.$ As an illustration, the 2G band corresponds to the frequency set $\mathcal{F}_{\mathrm{2G}}=\{935.0,945.0,955.0\}\,\text{MHz},$ so the aggregate 2G EMF can be computed as $E_{\mathrm{2G}}(t)=\sqrt{E^{2}_{935.0}+E^{2}_{945.0}+E^{2}_{955.0}}.$
Fig.  5(a) depicts the correlation of EMF among all frequencies. We note that the EMF levels across the bands (e.g., 700–900 MHz, 1800–2100 MHz, 2.6 GHz, and 3.4–3.8 GHz) tend to be more strongly correlated with each other indicating that exposure dynamics are largely driven by co-located carriers within the same spectral neighbourhood.
Fig. 5(b) depicts the correlation of EMF observed among the cellular technologies in the Italian dataset. 2G and 3G exhibit the strongest off-diagonal correlation ($\rho\approx 0.61$), reflecting similar temporal usage patterns of legacy technologies at this location. In contrast, 4G is only weakly correlated with the other technologies ($\rho\leq 0.19$), while 5G shows moderate correlation with 2G and 3G ($\rho\approx 0.27$$0.37$).
The EMF correlation map (Fig. 5(c)) reveals a high correlation among TIM, VF, and W3, with pairwise correlations in the range $\rho\approx 0.56$$0.65$, while Iliad remains only weakly correlated with the other operators ($\rho\approx 0.11$$0.15$). This indicates that the EMF exposure patterns of the three incumbent operators are highly synchronized at the hospital site, whereas Iliad follows a more independent deployment and load profile.
TABLE III: A comparative analysis of univariate forecasting performance of EMFusion with imputation without external conditions. Metrics are ordered as CRPS, ND, NRMSE, MAPE, and PICP. Lower is better for CRPS, ND, NRMSE, MAPE; higher is better for PICP. Bold = best in column; underline = second best in column.
Operator-wise Performance
Iliad TIM VF W3
Model CRPS ND NRMSE MAPE PICP CRPS ND NRMSE MAPE PICP CRPS ND NRMSE MAPE PICP CRPS ND NRMSE MAPE PICP
NF 0.0224 0.1901 0.2271 22.0056 69.36 0.0110 0.2674 0.3344 30.5366 65.99 0.0057 0.1505 0.1899 16.3322 69.26 0.0134 0.1732 0.2126 19.5926 72.44
IQLSTM 0.0208 0.1742 0.2088 22.0082 69.61 0.0098 0.2272 0.2905 23.7362 65.43 0.0060 0.1484 0.2158 16.5989 64.66 0.0146 0.1742 0.2179 19.3497 67.87
DDPM 0.0281 0.2266 0.2752 24.7179 75.38 0.0131 0.3078 0.4063 35.0737 76.78 0.0067 0.1741 0.2173 19.2858 90.22 0.0338 0.3945 0.4217 42.9547 28.02
VAE 0.0255 0.2040 0.2424 24.0345 54.86 0.0121 0.2880 0.3504 33.7195 53.63 0.0063 0.1630 0.2040 17.8781 58.19 0.0146 0.1842 0.2236 21.5268 66.11
LSTM-Dropout 0.0311 0.2298 0.2626 32.2884 33.07 0.0137 0.2944 0.3334 38.0923 31.38 0.0072 0.1687 0.2004 20.5388 31.19 0.0232 0.2515 0.2747 34.6254 31.17
WGAN 0.0237 0.1856 0.2216 21.6153 50.18 0.0207 0.3413 0.3809 44.7403 0.00 0.0055 0.1419 0.1780 15.6357 60.28 0.0138 0.1741 0.2096 20.5489 62.22
TimeGrad (29) 0.0329 0.3405 0.8942 89.2536 0.0094 0.4230 1.3310 52.4586 0.0084 0.2997 1.0983 29.7099 0.0203 0.3184 1.0497 48.2214
TimeGrad (8)v2 0.0552 0.1790 0.5465 26.7691 0.0271 0.3545 1.1546 34.5811 0.0201 0.1785 1.0200 15.6345 0.0367 0.2124 0.7484 18.9919
EMForecaster 0.0243 0.1706 0.2019 21.5334 34.50 0.0103 0.2042 0.2499 22.5550 29.67 0.0061 0.1419 0.2031 16.0613 44.94 0.0147 0.1585 0.1937 18.5124 37.61
EMFusion 0.0169 0.1304 0.1772 14.2795 50.40 0.0067 0.1433 0.2143 14.1847 58.51 0.0043 0.0997 0.1781 11.0273 52.25 0.0099 0.1132 0.1639 12.1131 52.19
Technology-band-wise Performance
2G 3G 4G 5G
Model CRPS ND NRMSE MAPE PICP CRPS ND NRMSE MAPE PICP CRPS ND NRMSE MAPE PICP CRPS ND NRMSE MAPE PICP
NF 0.0028 0.1416 0.1773 14.3473 59.79 0.0150 0.2153 0.2534 24.2409 62.59 0.0217 0.1637 0.1983 18.1657 69.23 0.0024 0.2103 0.3111 18.5137 72.20
IQLSTM 0.0027 0.1342 0.1708 13.7036 56.70 0.0134 0.1841 0.2256 20.9445 63.17 0.0210 0.1488 0.1827 17.1465 66.22 0.0028 0.1901 0.2994 18.1656 66.50
DDPM 0.0029 0.1462 0.1858 15.0905 82.02 0.0150 0.2132 0.2536 26.5512 76.83 0.0267 0.1890 0.2345 19.6537 74.21 0.0031 0.2536 0.3570 24.2072 71.12
VAE 0.0028 0.1429 0.1786 14.3644 60.92 0.0168 0.2227 0.2606 26.1413 45.59 0.0237 0.1698 0.2051 19.2985 58.84 0.0028 0.2298 0.3246 21.2987 57.09
LSTM-Dropout 0.0033 0.1388 0.1699 14.4200 32.48 0.0196 0.2728 0.2991 44.0706 48.00 0.0294 0.1994 0.2273 24.2642 39.79 0.0038 0.2742 0.3224 28.7500 14.29
WGAN 0.0028 0.1374 0.1747 13.7520 53.91 0.0146 0.2003 0.2339 23.7100 54.25 0.0224 0.1556 0.1903 17.8671 51.53 0.0046 0.2674 0.3303 26.9596 0.00
TimeGrad (29) 0.0059 0.3761 2.2558 38.7363 0.0267 0.3133 0.9559 61.6179 0.0190 0.3375 0.9451 56.1075 0.0051 0.3912 0.9879 40.6475
TimeGrad (8)v2 0.0072 0.1746 1.9840 16.9011 0.0339 0.2040 0.6425 24.7493 0.0584 0.1888 0.6344 21.9084 0.0111 0.3741 1.0243 21.8487
EMForecaster 0.0104 0.1682 0.2156 19.0429 37.40 0.0139 0.1688 0.2122 19.6655 36.68 0.0134 0.1661 0.2102 19.3257 37.34 0.0128 0.1802 0.2262 20.6762 34.69
EMFusion 0.0016 0.0821 0.1135 8.3460 59.64 0.0085 0.1153 0.1564 13.1791 52.88 0.0169 0.1109 0.1495 11.8418 55.69 0.0019 0.1518 0.4311 11.5743 66.95
Figure 6: Comparison of EMF Forecasts in Validation Set Across Technology Bands (Top) and Operators (Bottom)
TABLE IV: A comparative analysis of univariate forecasting performance of EMFusion with imputation without external conditions. Metrics are ordered as CRPS, ND, NRMSE, MAPE, and PICP. Lower is better for CRPS, ND, NRMSE, MAPE; higher is better for PICP. Bold = best in row; underline = second best in row.
Univariate Multivariate
Metric EMFusion (Uncond.) EMFusion (Seasons) EMFusion (Workingday) EMFusion (WorkingHour) EMFusion (multi) EMFusion (Uncond.) EMFusion (Seasons) EMFusion (Workingday) EMFusion (WorkingHour) EMFusion (multi)
Operator W3 CRPS 0.0099 0.0111 0.0103 0.0097 0.0193 0.1096 0.1095 0.1091 0.0086 0.0365
ND 0.1108 0.1265 0.1230 0.1156 0.2179 0.8993 0.8950 0.8935 0.0946 0.3300
NRMSE 0.1601 0.1708 0.1715 0.1669 0.2673 0.9086 0.9041 0.9028 0.1467 0.3960
MAPE 11.79 14.13 13.77 12.60 25.30 88.29 87.87 87.69 9.89 38.06
PICP 50.49 48.91 54.78 60.26 43.41 0.00 0.00 0.00 52.69 24.88
Iliad CRPS 0.0164 0.0173 0.0167 0.0153 0.0273 0.0724 0.0724 0.0680 0.0120 0.0435
ND 0.1215 0.1334 0.1319 0.1211 0.1975 0.4382 0.4382 0.4041 0.0903 0.2759
NRMSE 0.1744 0.1744 0.1795 0.1724 0.2380 0.4819 0.4819 0.4419 0.1469 0.3348
MAPE 14.05 15.94 15.05 13.93 26.39 40.38 40.38 38.15 10.40 31.89
PICP 50.15 52.66 53.51 56.40 39.08 11.37 11.37 10.76 63.02 27.86
TIM CRPS 0.0066 0.0072 0.0066 0.0063 0.0122 0.0475 0.0475 0.0482 0.0065 0.0215
ND 0.1381 0.1548 0.1522 0.1466 0.2572 0.7524 0.7524 0.7579 0.1374 0.3891
NRMSE 0.2094 0.2180 0.2254 0.2184 0.3219 0.7756 0.7756 0.7810 0.2126 0.4356
MAPE 13.73 16.20 15.91 15.02 25.63 72.80 72.80 73.31 13.81 43.49
PICP 57.05 51.81 57.11 57.68 39.58 0.03 0.03 0.00 54.97 25.31
VF CRPS 0.0043 0.0046 0.0044 0.0043 0.0077 0.0349 0.0349 0.0331 0.0036 0.0122
ND 0.1005 0.1118 0.1078 0.1043 0.1806 0.6308 0.6308 0.6061 0.0826 0.2498
NRMSE 0.1762 0.1878 0.1872 0.1862 0.2445 0.6438 0.6438 0.6210 0.1571 0.3039
MAPE 11.46 12.35 12.15 11.74 19.34 62.48 62.48 59.82 9.67 25.23
PICP 51.04 54.16 55.79 58.90 44.61 0.13 0.13 0.16 56.56 29.63
Technology Band 2G CRPS 0.0070 0.0076 0.0071 0.0068 0.0131 0.0640 0.0639 0.0635 0.0062 0.0234
ND 0.1165 0.1310 0.1276 0.1222 0.2186 0.7609 0.7594 0.7525 0.1049 0.2922
NRMSE 0.1819 0.1922 0.1947 0.1905 0.2779 0.7760 0.7745 0.7683 0.1721 0.3542
MAPE 12.33 14.22 13.94 13.12 23.42 74.52 74.38 73.61 11.13 29.57
PICP 52.86 51.62 55.90 58.95 42.53 0.05 0.05 0.05 54.74 29.04
3G CRPS 0.0093 0.0101 0.0095 0.0089 0.0166 0.0661 0.0661 0.0646 0.0077 0.0288
ND 0.1178 0.1316 0.1287 0.1219 0.2133 0.6802 0.6791 0.6654 0.1012 0.3200
NRMSE 0.1800 0.1877 0.1909 0.1860 0.2679 0.7025 0.7013 0.6867 0.1658 0.3805
MAPE 12.76 14.65 14.22 13.32 24.17 65.99 65.88 64.74 10.95 38.29
PICP 52.18 51.88 55.30 58.31 41.67 2.88 2.88 2.73 56.81 30.09
4G CRPS 0.0090 0.0097 0.0092 0.0086 0.0162 0.0670 0.0669 0.0655 0.0074 0.0274
ND 0.1160 0.1298 0.1268 0.1202 0.2113 0.6923 0.6911 0.6775 0.0994 0.2959
NRMSE 0.1783 0.1865 0.1893 0.1847 0.2662 0.7130 0.7117 0.6974 0.1638 0.3536
MAPE 12.59 14.45 14.04 13.16 23.90 67.33 67.21 66.03 10.78 32.58
PICP 51.98 51.83 55.30 58.49 42.01 2.48 2.48 2.35 56.50 28.18
5G CRPS 0.0085 0.0091 0.0086 0.0081 0.0149 0.0505 0.0505 0.0494 0.0071 0.0241
ND 0.1246 0.1387 0.1360 0.1296 0.2231 0.6435 0.6435 0.6315 0.1119 0.3413
NRMSE 0.1923 0.1995 0.2044 0.1989 0.2816 0.6692 0.6692 0.6562 0.1823 0.3899
MAPE 13.24 15.17 14.76 13.93 24.25 62.11 62.11 61.15 11.93 39.23
PICP 53.82 52.61 55.88 57.67 40.72 2.89 2.89 2.73 57.38 25.59
V Numerical Results and Discussions
This section presents a comprehensive evaluation of the proposed EMFusion model against state-of-the-art benchmarks. To assess forecasting performance, we define both deterministic (point) and probabilistic forecasts metrics followed by the key baselines and evaluation settings.
V-A Evaluation Metrics
Let $y_{i}$ denote the true observed value at time step $i$, $\hat{y}_{i}$ the corresponding median (p50) point forecast produced by the model, and $m$ the total number of samples in the test set. In what follows, we describe the point forecast and
V-A1 Point Forecast Metrics
evaluate the accuracy of a single predicted value (the median) against the true value. The considered metrics are listed as follows:
$\bullet$ Mean Absolute Percentage Error (MAPE): measures the average percentage difference between the predicted and actual values, and is defined as follows:
$\text{MAPE}=\frac{100}{m}\sum_{i=1}^{m}\frac{|y_{i}-\hat{y}_{i}|}{|y_{i}|}$ (28)
$\bullet$ Normalized Deviation (ND): ND metric measures the total absolute error relative to the total magnitude of the true values, is defined as follows:
$\text{ND}=\frac{\sum_{i=1}^{m}|y_{i}-\hat{y}_{i}|}{\sum_{i=1}^{m}|y_{i}|}$ (29)
$\bullet$ Root Mean Square Error (RMSE) and Normalized RMSE (NRMSE): RMSE measures the standard deviation of the prediction errors. NRMSE normalizes RMSE by the root mean square of the true values, facilitating comparison across datasets with different scales.
$\text{RMSE}=\sqrt{\frac{1}{m}\sum_{i=1}^{m}(y_{i}-\hat{y}_{i})^{2}}$ (30)
$\text{NRMSE}=\frac{\text{RMSE}}{\sqrt{\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}}}$ (31)
V-A2 Probabilistic Forecast Metrics
Probabilistic metrics evaluate the quality of the predicted intervals and the full distribution, assessing both sharpness (the width of the interval) and reliability (the coverage).
$\bullet$ Prediction Interval Coverage Probability (PICP): measures the percentage of true values that fall within the model’s predicted interval. Let the prediction interval at time step $i$ be defined by its lower and upper quantiles, $L_{i}=q_{\underline{\alpha}}^{\,i}$ and $U_{i}=q_{\overline{\alpha}}^{\,i}$, corresponding to a nominal confidence level $\gamma=1-\alpha$. For example, an 80% prediction interval corresponds to $\gamma=0.8$ and $\alpha=0.2$. We define the coverage indicator as follows:
$c_{i}=\begin{cases}1,&\text{if }y_{i}\in[L_{i},U_{i}],\\[4.0pt] 0,&\text{otherwise}.\end{cases}$ (32)
The PICP is then computed as follows:
$\mathrm{PICP}=\frac{100}{m}\sum_{i=1}^{m}c_{i}.$ (33)
Ideally, $\mathrm{PICP}$ should be close to the nominal value $100\gamma$, indicating well-calibrated predictive intervals.
$\bullet$ Continuous Ranked Probability Score (CRPS): evaluates the quality of the entire predictive distribution by measuring the discrepancy between the predictive CDF $F$ and the observed value $y$ as shown below:
$\mathrm{CRPS}(F,y)=\int_{-\infty}^{\infty}\bigl(F(z)-\mathbf{1}\{y\leq z\}\bigr)^{2}\,\mathrm{d}z,$ (34)
where $\mathbf{1}\{\cdot\}$ denotes the indicator function. When only predictive samples $X^{(1)},\ldots,X^{(S)}\sim F$ are available, we approximate $F$ using the empirical CDF, i.e.,
$\widehat{F}(z)=\frac{1}{S}\sum_{s=1}^{S}\mathbf{1}\{X^{(s)}\leq z\}$ (35)
and compute CRPS at each time step accordingly [RasulEtAl2021TimeGrad].
V-B Considered Baselines
To validate the performance of EMFusion, we conduct a comparative analysis against a suite of forecasting models. A brief description of each basline model is provided below:
  • DDPM [ref:diffusion-energy]: the model uses a U-Net denoiser with time-dependent noise embeddings and is trained to match the true forward diffusion process via noise-prediction.
  • WGAN [huang2025probabilistic]: the generator maps historical EMF trajectories and latent noise to future EMF scenarios, while the critic enforces distributional realism through the Wasserstein loss with gradient penalty.
  • VAE [ref:vae]: the encoder maps historical EMF sequences into a low-dimensional latent space, and the decoder reconstructs future EMF trajectories from sampled latent variables conditioned on the history. This yields a parametric predictive distribution via the latent Gaussian prior.
  • Normalizing Flow (NF) [ref:diffusion-energy]: Starting from a base distribution (e.g., multivariate Gaussian), a sequence of invertible transformations is learned to model the joint distribution of future EMF values given the past values, enabling exact likelihood evaluation and sampling.
  • Improved Quantile LSTM (IQLSTM): follows [ref:iqlstm], where an LSTM-based temporal encoder outputs multiple conditional quantiles of the future EMF distribution. The model is trained using a pinball (quantile) loss over a set of pre-defined quantile levels, providing a non-parametric representation of forecast uncertainty for each time step.
  • Dropout-based Probabilistic Model: Following the Monte Carlo Dropout approach of Gal and Ghahramani [ref:dropout], this algorithm employs a deep LSTM forecaster with dropout layers activated both during training and inference, byy running multiple stochastic forward passes with different dropout masks
  • TimeGrad: TimeGrad [RasulEtAl2021TimeGrad] is a diffusion-based time-series forecasting model that combines an autoregressive RNN backbone with a conditional diffusion head.
V-C Evaluation Settings
All models were trained using an identical training dataset. Their respective hyperparameters were optimized via a grid search procedure on the validation datasets. All algorithms were executed using PyTorch on the Trillium high-performance computing cluster, hosted by SciNet at the University of Toronto. Specifically, we utilized the GPU-accelerated nodes, which are equipped with an AMD EPYC 9654 (Zen 4) 2.4 GHz CPU, 749 GB of system RAM, and four NVIDIA H100 SXM GPUs (80 GB memory each).
For each day, the EMF time series contains $192$ regularly sampled observations. We use a 7-day lookback window, randomly sampled from the dataset, to predict the subsequent day. Formally, the historical input window is $H=7\times 192=1344$ time steps and the prediction window is $F=192$ steps. During sampling,we generate $\hat{N}=100$ scenarios for probabilistic forecasting. The hyperparameters for the EMFusion are set as follows: For training, the learning rate is 5e45e-4, the batch size is 64, and the model is trained for $1500$ epochs. The U-Net model has a depth of 6, and the attention mechanism uses $8$ heads. In the simulations, physical constants include the free-space wave impedance $Z_{0}=376.73\ \Omega$ and the antenna-gain calibration constant $A_{GC}=9.73$.
V-D Performance Evaluation
V-D1 Comparative Analysis of EMFusion without External Conditions
Table-III analyzes the univariate forecasting performance of the EMFusion model, configured with imputation and without external conditions. We compare it against several baselines, including generative models (NF, DDPM, VAE, WGAN) and other probabilistic forecasting models (IQLSTM, LSTM-Dropout). The evaluation is based on five key metrics, CRPS, ND, NRMSE, and MAPE (where lower is better) and PICP (where higher is better).
In both the operator-wise and technology-band-wise breakdowns, EMFusion consistently achieves the best scores for CRPS, ND, NRMSE, and MAPE, while only WGAN achieves a fractionally better NRMSE for the VF dataset. From the operator perspective, the performance gap is even more pronounced. EMFusion’s CRPS (0.0067) and MAPE (14.18%) are far superior to all baselines in terms of TIM. We also observe EMFusion drastically reduces all error metrics in the 2G 3G and 4G dataset. While 5G dataset appears to be the most challenging, with higher error values for all models. Despite this, EMFusion still delivers the best performance in CRPS (0.0019), ND (0.1518), and MAPE (11.57%).
In terms of evaluation, the highest PICP scores are consistently achieved by DDPM and NF. For example, DDPM achieves an exceptional 90.22% for VF and 82.02% for 2G, while NF scores 72.44% for W3. This indicates that these models enable larger percentage of the true (observed) values fall inside the model’s prediction intervals. EMFusion provides a more balanced result, with PICP scores generally in the 50-60% range. This suggests it produces tighter, more accurate prediction intervals that are less conservative but more closely reflect the actual forecast.
Fig. 6 highlights the performance of multivariate EMFusion across all datasets organized by cellular technology bands and operators. Multivariate EMFusion with working hour condition provides the closest match to the ground truth EMF values, i.e., it reproduces the daily exposure cycle, aligns well with the timing of peaks and troughs, and preserves the moderate intra-day variability without following high-frequency noise. Univariate EMFusion with working hour condition exhibits a similar pattern but tends to slightly underestimate the most pronounced peaks in some panels. In contrast, NF and VAE generally produce smoother forecasts that track the overall trend but miss sharp local excursions, leading to significant phase and amplitude mismatches. DDPM is more responsive to transient fluctuations, but this often manifests as spurious oscillations and overfitting to local noise, especially in higher-exposure bands such as 4G and 5G.
TABLE V: Performance comparison Lower is better for CRPS, ND, NRMSE, MAE, MAPE, and Energy Score; for PICP, a value closest to 80% is preferred. Best per column in bold. Reported as mean $\pm$ standard error.
Method CRPS ND NRMSE MAE MAPE PICP Energy Score
NF 0.0110 $\pm$ 0.0070 0.1739 $\pm$ 0.0332 0.2447 $\pm$ 0.0941 0.0337 $\pm$ 0.0219 18.1875 $\pm$ 2.8910 69.1871 $\pm$ 6.0702 0.0163 $\pm$ 0.0103
IQLSTM 0.0120 $\pm$ 0.0076 0.1829 $\pm$ 0.0338 0.2575 $\pm$ 0.0954 0.0351 $\pm$ 0.0226 19.2144 $\pm$ 2.6370 61.8676 $\pm$ 5.9028 0.0192 $\pm$ 0.0120
WGAN 0.0123 $\pm$ 0.0071 0.1959 $\pm$ 0.0507 0.2588 $\pm$ 0.0998 0.0356 $\pm$ 0.0210 21.3099 $\pm$ 4.3154 50.0930 $\pm$ 9.8127 0.0202 $\pm$ 0.0114
VAE 0.0128 $\pm$ 0.0083 0.1908 $\pm$ 0.0411 0.2597 $\pm$ 0.0971 0.0367 $\pm$ 0.0237 20.6526 $\pm$ 3.8411 52.2893 $\pm$ 4.4282 0.0213 $\pm$ 0.0139
DDPM 0.0132 $\pm$ 0.0085 0.1989 $\pm$ 0.0520 0.2647 $\pm$ 0.0985 0.0385 $\pm$ 0.0260 22.5219 $\pm$ 6.0977 85.9521 $\pm$ 6.4352 0.0156 $\pm$ 0.0108
Dropout 0.0179 $\pm$ 0.0117 0.2418 $\pm$ 0.0481 0.2937 $\pm$ 0.0867 0.0470 $\pm$ 0.0299 30.2881 $\pm$ 6.8275 27.9310 $\pm$ 6.7606 0.0328 $\pm$ 0.0218
EMFusion-Uncond. 0.0212 $\pm$ 0.0154 0.2486 $\pm$ 0.0688 0.3257 $\pm$ 0.1048 0.0523 $\pm$ 0.0377 23.6121 $\pm$ 6.2094 33.0397 $\pm$ 10.4381 0.0389 $\pm$ 0.0284
EMFusion-MV(Multi) 0.0154 $\pm$ 0.0102 0.2083 $\pm$ 0.0412 0.2808 $\pm$ 0.0955 0.0411 $\pm$ 0.0266 22.8826 $\pm$ 5.1239 41.6521 $\pm$ 6.5349 0.0275 $\pm$ 0.0185
EMFusion-MV(Working Day) 0.0696 $\pm$ 0.0634 0.6345 $\pm$ 0.2315 0.6705 $\pm$ 0.2088 0.1433 $\pm$ 0.1262 62.0675 $\pm$ 22.9755 4.8629 $\pm$ 7.7988 0.1381 $\pm$ 0.1270
EMFusion-MV(Season) 0.0700 $\pm$ 0.0635 0.6355 $\pm$ 0.2347 0.6735 $\pm$ 0.2123 0.1445 $\pm$ 0.1265 61.9106 $\pm$ 23.5813 5.7850 $\pm$ 8.1614 0.1388 $\pm$ 0.1272
EMFusion-MV(WorkingHour) 0.0071 $\pm$ 0.0043 0.1049 $\pm$ 0.0241 0.1858 $\pm$ 0.0936 0.0195 $\pm$ 0.0120 10.7561 $\pm$ 1.4898 56.4799 $\pm$ 5.0838 0.0118 $\pm$ 0.0072
V-D2 Univariate vs Multivariate Conditional EMFusion
Table-IV shows the significance of incorporating appropriate contextual conditions in the reverse diffusion process. In particular, the WorkingHour-conditioned EMFusion achieves the lowest CRPS, ND, NRMSE, and MAPE regardless of the operator (W3, Iliad, TIM, VF) or technology band (2G–5G). The multi-conditioned multivariate model (last column) does not surpass the WorkingHour-only model, suggesting that fine-grained time-of-day conditioning captures most of the useful temporal structure for EMF exposure, while additional context yields diminishing returns. From the operator perspective, TIM and VF benefit strongly from conditional diffusion: for example, under multivariate modeling their CRPS and MAPE values under EMFusion (WorkingHour) are an order of magnitude smaller than uncondition EMFusion, indicating substantially sharper and more accurate forecasts.
From the technology-band viewpoint, EMFusion sharply reduces all error metrics for legacy bands (2G, 3G, 4G), with the largest relative gains observed for 4G, where EMFusion (WorkingHour) approximately halves the CRPS and MAPE compared to unconditioned EMFusion in the multivariate setting. The 5G band remains the most challenging case, with uniformly higher errors across all models due to its more volatile traffic and deployment patterns. Nonetheless, EMFusion still delivers the best multivariate performance in CRPS, ND, and MAPE for 5G, indicating that the diffusion-based approach remains robust even under rapidly fluctuating EMF dynamics. In terms of interval quality, unconditioned EMFusion exhibits very low PICP values in the multivariate case, reflecting overly narrow and under-dispersed prediction intervals. EMFusion, by contrast, achieves PICP values predominantly in the 50-65% range across operators and bands, with the WorkingHour configuration again performing best. Although these values are below the ideal 80% target, they indicate tighter, less conservative intervals that are more informative than those from unconditioned EMFusion.
Table-V reports the forecasting performance of all methods, averaged over all carriers and technologies. The proposed EMFusion-MV-WorkingHour consistently outperforms across all error- and distribution-based scores, indicating sharper and better-calibrated predictive distributions. Although its PICP is slightly lower than that of DDPM, it remains competitive with the strongest baselines while delivering substantially improved overall forecasting quality.
VI Conclusion
Existing EMF forecasting approaches do not fully exploit the spatio-temporal dependencies across multiple frequency bands, operators, and contextual conditions. This study introduced EMFusion, a conditional diffusion based probabilistic forecasting framework. The framework employs conditional embeddings, cross-attention based context integration, and imputation-based sampling to improve temporal coherence. Experiments on real EMF datasets showed that EMFusion achieved lower deterministic forecasting errors and better probabilistic calibration than representative deep learning time-series and generative baselines across different operators and frequency selections, and that the predicted point values and prediction intervals were closer to the observed EMF trajectories. Future work may incorporate base-station location, transmit power, user distribution, and reconfigurable intelligent surface deployments as additional conditioning variables to further tighten the forecast distribution.
Acknowledgment
The authors would like to thank Daniele Franci and Settimio Pavoncello for providing the EMF measurement hardware used to collect the Italian dataset.